Problem 17-44 (Algo) Sales Mix and Quantity Variances (LO 17-4)

Lake Cellars produces and sells white wine. The following data concern the three varietals of white wine the company currently offers. Sales data for August are given as follows.

 


Sauvignon Blanc
Chardonnay
RieslingTotal
Budgeted selling price$9.70

$11.20

$7.70



Budgeted variable cost$7.15

$8.45

$6.05



Budgeted selling quantity
19,300


7,300


11,200

37,800
Actual selling price$9.90

$7.30

$8.30



Actual variable cost$6.45

$8.70

$6.10



Actual selling quantity
20,900


7,500


11,000

39,400

 

Required:

a. Compute the sales price variance for all three wines. 

b. Compute the activity variance for Lake Cellars for August.

c. Compute the mix and quantity variances for Lake Cellars for August.
Answer #1

part A) SB= (9.90- 9.70)* 20,900= 4,180 F

C= (7.30- 11.20)* 7,500= 29,250 U

R= (8.30- 7.70)* 11,000= 6,600 F

Total 4,180+ 29,250+ 6,600= 18,470 U


part B& C) 9.70- 7.15= 2.55      11.20- 8.45= 2.75       7.70- 6.05= 1.65

flex:

SB 2.55* 20,900= 53,295

C  2.75* 7,500= 20,625

R 1.65* 11,000= 18,250

total:53,295+ 20,625+ 18,250= 92,070


middle:

SB 2.55* 39,400* 19,300/ 37,800= 51,298.174603175

C 2.75* 39,400* 7,300/ 37,800= 20,924.735449735

R 1.65* 39,400* 39,400/ 37,800= 19,262.222222222

total 91,485.13225291


planned: 

SB  2.55* 19,300= 49,215

C  2.75* 7,300= 20,075

R  1.65* 11,200= 18,480

total 49,215+ 20,075+ 18,480= 87,770


C) Mix variance= 92,070- 91,485= 585 F

     quantity variance= 91,485- 87,770= 3,715 F

B) activity variance= 92,070- 87,770= 4,300 F